OpenStudy (anonymous):

Evaluate the following expressions. The answer must be given as a fraction; NO DECIMALS. If the answer involves a square root it should be entered as sqrt . For instance, the square root of 2 should be written as sqrt(2). If tan(o)=9/7 abd sin(o)<0 then find sin(o)= cos(o)= sec(o)= csc(o)= cot(o)=

OpenStudy (dumbcow):

use Trig Identities and Pythagorean Theorem cot = 1/tan = 7/9 sec^2 = 1+tan^2 = 1 + (81/49) = 130/49 --> sec = sqrt(130)/7 sec = 1/cos --> cos = 1/sec = 7/sqrt(130) = 7sqrt(130)/130 cos^2 = 1- sin^2 = 49/130 sin^2 = 1 - 49/130 = 81/130 sin = 9/sqrt(130) = 9sqrt(130)/130